# Language-based Audio Moment Retrieval

### v1

@@ Abstract
In this paper, we propose and design a new task called audio moment retrieval (AMR). Unlike conventional language-based audio retrieval tasks that search for short audio clips from an audio database, AMR aims to predict relevant moments in untrimmed long audio based on a text query. Given the lack of prior work in AMR, we first build a dedicated dataset, Clotho-Moment, consisting of large-scale simulated audio recordings with moment annotations. We then propose a DETR-based model, named Audio Moment DETR (AM-DETR), as a fundamental framework for AMR tasks. This model captures temporal dependencies within audio features, inspired by similar video moment retrieval tasks, thus surpassing conventional clip-level audio retrieval methods. Additionally, we provide manually annotated datasets to properly measure the effectiveness and robustness of our methods on real data. Experimental results show that AM-DETR, trained with Clotho-Moment, outperforms a baseline model that applies a clip-level audio retrieval method with a sliding window on all metrics, particularly improving Recall1@0.7 by 9.00 points. Our datasets and code are publicly available in https://h-munakata.github.io/Language-based-Audio-Moment-Retrieval.
@@ I Introduction
Language-based audio retrieval, referred to simply as audio retrieval, is to search for desired audio from an audio database using a natural language query. Researchers pay attention to this technology because of its wide range of applications, such as a retrieval system of historical sound archives and sound effects [1]. With an introduction of large-scale audio-text datasets [2, 3, 4], various approaches have been proposed [5, 6, 7, 8, 9].
@@
Although existing audio retrieval models assume that the audio is trimmed as a short clip, there is a great demand for retrieving specific time segments from untrimmed long audio. Potential applications include, for example, highlight moment detection for sports broadcasts and criminal moment detection for surveillance systems. These applications motivate us to retrieve moments in the audio using natural language queries. Figure 1 shows the retrieval system we envision to implement the application. Given a query “Spectators watch sports and cheer” along with long audio, the model is expected to retrieve audio moments as start and end timestamps (16 to 44 seconds). A straightforward solution to building such moment retrievers is to utilize existing audio retrieval methods in three steps: dividing the long audio into multiple clips, forwarding them into the audio retrieval models, and calculating the similarity between the text query and the audio clip embeddings in the shared embedding space. However, this approach is sub-optimal because it processes the audio clip sequence independently, failing to capture the temporal dependencies between the clips.
Figure 1: An overview of audio moment retrieval (AMR). Given a long audio and text query, the system retrieves the relevant moments as the start and end time stamps.
To address this limitation, we can draw inspiration from the field of the video domain, where modeling temporal dependencies is well explored. Video moment retrieval (VMR) is a similar task that retrieves specific moments from a long video [11, 12, 13, 14, 15, 16, 17, 18, 19]. For this task, DEtection TRansformer (DETR)-based models [20] show high retrieval performance [15, 17, 18]. These models learn not only temporal dependencies inside the video frames but also cross-modal similarity between frames and text query, resulting in correct moment prediction. Various methods have been proposed in VMR and we think the ideas of these methods are helpful to our target system.
Figure 2: The left figure describes Clotho-Moment. By overlaying Clotho on Walking Tour, audio and text query-audio moment pairs are generated. The right side of the figure describes the architecture of the proposed Audio Moment DETR. This model first extracts embedding by audio/text encoder and then the DETR-based network transforms the embedding to pairs of the audio moment and its confidence score.
@@ II Difference to Existing Tasks
AMR has three similar tasks in audio and vision domains: sound event detection (SED), target source extraction (TSE), and video moment retrieval (VMR). In this section, we emphasize the novelty of AMR by enumerating the difference between AMR and these tasks.
Sound event detection (SED) is the task of predicting both the sound event class labels and the event time boundaries from the input audio [21, 22, 23, 24, 25, 26]. SED assumes the target labels are predetermined and each label corresponds to a single event. In contrast, AMR assumes an open vocabulary for queries and thus can handle queries related to multiple events. Hence, AMR can be seen as zero-shot SED in an open-vocabulary setting. In fact, we experimented with zero-shot SED using the AMR model (Section V-C).
Target source extraction (TSE) is the task of extracting sources relevant to the text query from the input mixture [27, 28]. The difference between AMR and TSE lies in the target output. AMR aims to identify the relevant moment, while TSE aims extract the relevant separated source signal. Therefore, even if we obtain separated signals with TSE, we still need to perform AMR to retrieve the desired moments.
@@ III Dataset for Audio Moment Retrieval
We construct datasets to retrieve audio moments relevant to the text query from untrimmed audio data. The audio we assume for AMR is one minute long and includes some specific scenes that can be represented by text such as audio of a sports broadcast including a scene in which the goal is scored and people cheering. Based on these assumptions, we create two types of datasets.
@@ III-A Clotho-Moment
We propose Clotho-Moment, a large-scale simulated AMR dataset. The dataset is generated based on the simulation by leveraging existing audio-text data, which does not require additional annotation. We generated this dataset from two datasets, Clotho [2] and WalkingTour [29]. Clotho consists of various manually annotated 15 to 30-second audio clip data collected on Freesound. Each audio sample has five captions comprising 8 to 20 words. This dataset has development, validation, and evaluation splits and each split contains 3839, 1045, and 1045 audio samples. Walking Tour contains ten videos over an hour long recorded in various cities. We use only the audio in the video. Following Clotho, we split this dataset into development, validation, and evaluation splits in a 7:1:2 ratio.
Clotho-Moment simulates long recorded audio in the city including some scenes contained in Clotho. As shown on the left side of Figure 2, this dataset is generated with Clotho and Walking tour audio samples as foreground and background, respectively. Figure 2. Inspired by the simulation method proposed for the speaker diarization [30], the audio and pairs of text query and moments are generated by overlaying Clotho samples on Walking Tour samples at randomly sampled intervals as shown in Algorithm 1. Note that our method can only generate a single audio moment per query and each scene is not overlapped.
We make several adjustments to make the data more realistic. To simulate scenes with various volume levels, the signal power of the sampled audio of Clotho and Walking Tour is randomly weighted at ±$\pm 5$ dB, and at $-20\pm 5$ dB, respectively. To make audio moments more accurate, we removed the silent onsets and offsets of Clotho clips that were 20 dB lower than the overall signal power. We set hyperparameter $\lambda$ controlling the average interval length to $30$ second. To reduce mismatch with our manually annotated dataset created described below, Walking Tour samples are divided into one-minute segments in one-second intervals in advance. Finally, we generated 32694, 4918, and 6649 samples for development, validation, and evaluation split.
@@
0:   $\mathcal{D}_{\rm Clotho}$ // Set of of query-foreground audio pairs      ${\mathcal{D}_{\rm WalkingTour}}$ // Set of background audio
0:   $\mathbf{x}$ // Generated audio        $Y=\{(q_{1},t_{\rm 1,start},t_{\rm 1,end}),...\}$ // Set of query-moment pairs
1:  $Y\leftarrow\emptyset,~{}t=0$ // Initialize
2:  $\mathbf{x}\sim\mathcal{D}_{\rm WalkingTour}$ // Sample background audio
@@
8:     if $t_{\rm end}>{\rm duration}(\mathbf{x})$ then
9:        break
10:     end if
11:     $\mathbf{x}[t_{\rm start}:t_{\rm end}]=\mathbf{x}[t_{\rm start}:t_{\rm end}]+\mathbf{x}_{\rm moment}$ // Overlay
12:     $Y\leftarrow Y~{}|~{}\{(q,t_{\rm start},t_{\rm end})\}$ // Update sets
13:     $t\leftarrow t_{\rm end}$
14:  end while
15:  return  $\mathbf{x},Y$
@@ III-B Real Data for Evaluation
In addition to the simulated data, we create a small-scale manually annotated AMR dataset to evaluate the retrieval performance in more realistic environments. As a relatively long video data collection containing various scenes, we chose UnAV-100, an untrimmed YouTube video dataset [31], and extracted audio from the video. We selected 77 videos and annotated 100 new text queries and audio moments. To make the evaluation easy, each query corresponds to a single audio moment. The audio duration is from 45 to 60 seconds and the resolution of the moments is one second. The captions consist of 6.3 words on average.
@@ IV Audio moment retrieval
In this section, we define audio moment retrieval (AMR). To tackle AMR, we propose Audio Moment DETR (AM-DETR) inspired by the VMR models [15, 17, 18]. By training with Clotho-Moment, AM-DETR enables to capture of both temporal dependencies inside the audio features and cross-modal similarity between audio features and text query.
@@ IV-A Task Formulation
We define audio moment retrieval as the task of predicting audio moments corresponding to a given text query. Let the input audio be denoted as $\mathbf{x}$, the text query as $q$, and the output audio moments as $y$. Here, $y$ represents the audio moment relevant to the query and may consist of multiple moments: $y=(y_{1},y_{2},\ldots,y_{N})$. Each moment is represented as a tuple of start and end times, $y_{n}=(t_{n,\text{start}},t_{n,\text{end}})$. This task can be viewed as a mapping problem where, given the input audio $\mathbf{x}$ and the text query $q$, we aim to find the relevant audio moments $y$. Formally, this can be expressed as $y=f_{\text{AMR}}(\mathbf{x},q)$. The key challenge is how to model the function $f_{\text{AMR}}$.
@@ IV-B Audio Moment DETR
We propose Audio Moment DETR (AM-DETR), an AMR model inspired by the model proposed for VMR [15, 17, 18]. As shown on the right side of Figure 2, AM-DETR consists of a feature extractor and a DETR-based network the same as the VMR models. The key part of AM-DETR is the DETR-based network [15, 17, 18] that captures the temporal dependency between audio features. Specifically, the architecture of AM-DETR is based on QD-DETR, which is simple but achieves high retrieval performance for VMR by capturing both the temporal dependency between video frames and cross-modal dependency between frames and text query [17].
@@
Finally, Transformer decoder transforms $\mathbf{z}^{\prime}$ to $K$ candidate audio moments $\hat{y}_{1},...,\hat{y}_{K}$ and their confidence scores $c_{1},...,c_{K}\in[0,1]$. The confidence score represents how plausible the corresponding predicted moment is. Here, $K$ is the hyperparameter controlling the number of outputs that assumes sufficiently larger than the number of the ground truth moment. To output $K$ candidates, the decoder also takes $K$ trainable anchors in addition to $\mathbf{z}^{\prime}$. Finally, the decoder is formulated as follows:
@@
Here, it is empirically known that using $c_{\phi(n)}$ not but $\log c_{\phi(n)}$ improves the performance in Eq (9[20]. We denote the optimal index as $\tilde{\phi}=\underset{\phi\in\mathcal{P}}{\mathrm{argmin}}~{}\mathcal{L}_{\rm match}(\hat{y},y,c,\phi)$ where $\mathcal{P}$ is a set of every $K!$ permutation. Finally, the overall loss is obtained by the weighted sum of the moment loss and the CE loss as follows:
@@
Method Feature extractor Win len Clotho-Moment (eval) UnAV-100 subset TUT-Sound Event 2017
@@ V-A Model Configuration
Proposed model. We trained AM-DETR using Clotho-Moment. To investigate the impact of the feature extractor, We used three feature extractors, CLAP [5], VR, and VR without contrastive learning. Here, VR is an audio retrieval model consisting of VAST [34] and RoBERTa [35], which is trained with our large-scale dataset and significantly outperforms CLAP in the Clotho benchmark [36]. For comparison, we used the sliding window with 1, 4, and 7-second window lengths. The hyperparameter settings of the DETR-based network followed [17]. Both Transformer encoder and decoder consisted of two stacks of attention layers with 256 units and eight heads, and the dimension of the feed-forward layers was set to 1024. We used AdamW optimizer [37] with a learning rate of $1\times 10^{-4}$ and trained the models for 100 epochs with a batch size of 32.
Baseline. We used a model combining the existing audio retrieval model with the sliding window as a baseline. To obtain audio moments, the baseline measures the similarities between audio and text embeddings for each window frame and then binarizes the similarities by comparing them to a threshold. The confidence score is defined as the average similarity within the moment. To improve performance, we applied a median filter to the resulting binary sequence. The threshold for the binarization and the length of median filters were tuned with the validation split of Clotho-Moment.
@@ V-B Evaluation
For our evaluation, we used Clotho-Moment evaluation split, UnAV-100 subset. In addition to this, we also used TUT sound events 2017 (TUT) [38] used in SED for reference. Since the labels of TUT consist of too short words for the text query of AMR, We added prompts with an average of 4.8 words to each label.
@@ V-C Results
Proposed vs Baseline. AM-DETR significantly outperformed the baseline in all metrics as shown in Table I (rows 3 and 6). Notably, the improvement of R1@0.7 and average mAP for UnAV-100 subset was 9.00 and 11.82 points. This result suggests that capturing the temporal dependency between audio features is important for AMR.
Feature extractor. We investigate the impact of contrastive learning on the feature extractor. First, we found that contrastive learning of the feature extractor significantly improves the retrieval performance (rows 4 and 5). Notably, the improvement of R1@0.7 and average mAP for UnAV-100 subset was 29.00 and 23.11 points, respectively. Next, we investigate the impact of the architecture of the feature extractors. Comparing VR and CLAP (rows 5 and 6), the performance was comparable indicating that performance in AMR does not necessarily match performance in the conventional audio retrieval.
@@
AMR as zero-shot SED. Because the AMR models can be seen as zero-shot SED as explained in Section II, we evaluated the performance for SED as the frame-level classification. We performed zero-shot SED using event label names with prompts as the text query, and measured their performances based on conventional frame-level metrics of the SED task. We evaluated AM-DETR and the model that simply combines CLAP and the sliding window, we called it SED-CLAP. The thresholds were tuned with the development set to maximize the F1 score, and the threshold was set to 0.5 if it was not tuned. As a result, we found two crucial results. First, the AMR model achieved the F1 score of over 77% of the conventional SED model based on supervised learning, despite a zero-shot condition. Second, comparing SED-CLAP, AM-DETR is less sensitive to the threshold, which reduces tuning effort.
@@
TABLE II: Frame-level classification performance on TUT. All metrics are measured in micro average. The window lengths of SED-CLAP and AM-DETR were 7 and 1, respectively.

### v3

@@ Abstract
In this paper, we propose and design a new task called audio moment retrieval (AMR). Unlike conventional language-based audio retrieval tasks that search for short audio clips from an audio database, AMR aims to predict relevant moments in untrimmed long audio based on a text query. Given the lack of prior work in AMR, we first build a dedicated dataset, Clotho-Moment, consisting of large-scale simulated audio recordings with moment annotations. We then propose a Detection Transformer-based model, named Audio Moment DETR (AM-DETR), as a fundamental framework for AMR tasks. This model captures temporal dependencies within audio features, inspired by similar video moment retrieval tasks, thus surpassing conventional clip-level audio retrieval methods. Additionally, we provide manually annotated datasets to properly measure the effectiveness and robustness of our methods on real data. Experimental results show that AM-DETR, trained with Clotho-Moment, outperforms a baseline model that applies a clip-level audio retrieval method with a sliding window on all metrics, particularly improving Recall1@0.7 by 9.00 points. Our datasets and code are publicly available in https://h-munakata.github.io/Language-based-Audio-Moment-Retrieval.
@@ I Introduction
Language-based audio retrieval, referred to simply as audio retrieval, is to search for desired audio from an audio database using a natural language query. Researchers pay attention to this technology because of its wide range of applications, such as a retrieval system of sound archives and sound effects [1]. With an introduction of large-scale audio-text datasets [2, 3, 4], audio retrieval methods that process audio and text in an embedding space have been investigated [5, 6, 7, 8, 9].
@@
Although existing audio retrieval models assume that the audio is trimmed as a short clip, there is a great demand for retrieving specific time segments from untrimmed long audio. Potential applications include, for example, highlight moment detection for sports broadcasts and criminal moment detection for surveillance systems. These applications motivate us to retrieve moments in the audio using natural language queries. Figure 1 shows the retrieval system we envision to implement the application. Given a query “Spectators are cheering and shouting in a sports game” along with long audio, the model is expected to retrieve audio moments as start and end timestamps (16 to 44 seconds). A straightforward solution to building such moment retrievers is to utilize existing audio retrieval methods in three steps: dividing the long audio into multiple clips, forwarding them into the audio retrieval models, and calculating the similarity between the text query and the audio clip embeddings in the shared embedding space. However, this approach is sub-optimal because it processes the audio clip sequence independently, failing to capture the temporal dependencies between the clips.
Figure 1: An overview of audio moment retrieval. Given a long audio and text query, the system retrieves relevant moments as start and end time stamps.
To address this limitation, we can draw inspiration from the field of the video domain, where modeling temporal dependencies is well explored. Video moment retrieval (VMR) is a similar task that retrieves specific moments from a long video [11, 12, 13, 14, 15, 16, 17]. For this task, DEtection TRansformer (DETR)-based models [18] show high retrieval performance [14, 15, 16]. These models learn not only temporal dependencies inside the video frames but also cross-modal similarity between frames and text query, resulting in correct moment prediction. Various methods have been proposed in VMR and we think the ideas of these methods are helpful to our target system.
Figure 2: The left figure describes Clotho-Moment. By overlaying Clotho on Walking Tours, audio and text query-audio moment pairs are generated. The right side of the figure describes the architecture of the proposed Audio Moment DETR. This model first extracts embedding by audio/text encoder and then the DETR-based network transforms the embedding to pairs of the audio moment and its confidence score.
@@ II Difference to Existing Tasks
AMR has three similar tasks in audio and vision domains: sound event detection (SED), target sound extraction (TSE), and video moment retrieval (VMR). In this section, we emphasize the novelty of AMR by enumerating the difference between AMR and these tasks.
Sound event detection (SED) is the task of predicting both the sound event class labels and the event time boundaries from the input audio [19, 20, 21, 22, 23, 24, 25, 26, 27]. SED assumes the target labels are predetermined and each label corresponds to a single event. In contrast, AMR assumes an open vocabulary for queries and thus can handle queries related to multiple events. Hence, AMR can be seen as zero-shot SED in an open-vocabulary setting.
Target sound extraction (TSE) is the task of extracting sound sources relevant to the text query from the input mixture [28, 29]. The difference between AMR and TSE lies in the target output. AMR aims to identify the relevant moment, while TSE aims extract the relevant separated source signal. Therefore, even if we obtain separated signals with TSE, we still need to perform AMR to retrieve the desired moments.
@@ III Dataset for Audio Moment Retrieval
We construct datasets to retrieve audio moments relevant to the text query from untrimmed audio data. The audio we assume for AMR is one minute long and includes some specific scenes that can be represented by text, such as audio of a sports broadcast including a scene in which the goal is scored and people cheering. Based on these assumptions, we create two types of datasets.
@@ III-A Clotho-Moment
We propose Clotho-Moment, a large-scale simulated AMR dataset. The dataset is generated based on the simulation by leveraging existing audio-text data, which does not require additional annotation. Clotho-Moment is generated from two datasets, Clotho [2] and Walking Tours [30]. These datasets are selected because Clotho samples contain specific scenes that we assume to be the target of AMR, and most parts of Walking Tours samples do not. Clotho consists of various manually annotated 15 to 30-second audio clip data collected on Freesound. Each audio sample has five captions comprising 8 to 20 words. This dataset has development, validation, and evaluation splits and each split contains 3839, 1045, and 1045 audio samples. Walking Tours contains ten videos over an hour long recorded in various cities. We use only the audio in the video. Following Clotho, we split this dataset into development, validation, and evaluation splits in a 7:1:2 ratio.
Clotho-Moment simulates long recorded audio in the city including some scenes contained in Clotho. As shown on the left side of Figure 2, this dataset is generated with Clotho and Walking Tours audio samples as foreground and background, respectively. Figure 2. Inspired by the simulation method proposed for the speaker diarization [31], the audio and pairs of text query and moments are generated by overlaying Clotho samples on Walking Tours samples at randomly sampled intervals as shown in Algorithm 1. Note that our method can only generate a single audio moment per query and each scene is not overlapped.
We make several adjustments to make the data more realistic. To simulate scenes with various volume levels, the signal power of the sampled audio of Clotho and Walking Tours is randomly weighted at ±5$\pm$ 5± 5 dB, and at 20±5-20$\pm$ 5- 20 ± 5 dB, respectively. To make audio moments more accurate, we removed the silent onsets and offsets of Clotho clips that were 20 dB lower than the overall signal power. We set hyperparameter $\lambda$ controlling the average interval length to 303030 second. To reduce mismatch with our manually annotated dataset created described below, Walking Tours samples are divided into one-minute segments in one-second intervals in advance. Finally, we generated 32694, 4918, and 6649 samples for development, validation, and evaluation split.
@@
[…]
@@
8:  if $t_{\rm end}>{\rm duration}(\mathbf{x})$ then break
9:  $\mathbf{x}[t_{\rm start}:t_{\rm end}]=\mathbf{x}[t_{\rm start}:t_{\rm end}]+\mathbf{x}_{\rm moment}$ // Overlay
10:  $Y\leftarrow Y~|~\{(q,t_{\rm start},t_{\rm end})\}$ // Update sets
11:  $t\leftarrow t_{\rm end}$
12:end while
13:return  $\mathbf{x},Y$
@@ III-B Real Data for Evaluation
We also create a small-scale manually annotated AMR dataset to evaluate the retrieval performance in more realistic environments. As a relatively long video data collection containing various scenes, we chose UnAV-100, an untrimmed YouTube video dataset [32], and extracted audio from the video. We selected 77 videos and annotated 100 new text queries and audio moments. To make the evaluation easy, each query corresponds to a single audio moment. The audio duration is from 45 to 60 seconds and the resolution of the moments is one second. The captions consist of 6.3 words on average.
@@ IV Audio moment retrieval
In this section, we define audio moment retrieval (AMR). To tackle AMR, we propose Audio Moment DETR (AM-DETR) inspired by the VMR models [14, 15, 16]. By training with Clotho-Moment, AM-DETR enables to capture of both temporal dependencies inside the audio features and cross-modal similarity between audio and text.
@@ IV-A Task Formulation
We define audio moment retrieval as the task of predicting audio moments corresponding to a given text query. Let the input audio be denoted as $\mathbf{x}$, the text query as $q$, and the output audio moments as $y$. Here, $y$ represents the audio moment relevant to the query and may consist of multiple moments: $y=(y_{1},y_{2},\ldots,y_{N})$. Each moment is represented as a tuple of start and end times, $y_{n}=(t_{n,\text{start}},t_{n,\text{end}})$. This task can be viewed as a mapping problem where, given the input audio $\mathbf{x}$ and the text query $q$, we aim to find the relevant audio moments $y$. Formally, this can be expressed as $y\simeq f_{\text{AMR}}(\mathbf{x},q)$. The key challenge is how to model the function $f_{\text{AMR}}$.
@@ IV-B Audio Moment DETR
We propose Audio Moment DETR (AM-DETR), an AMR model inspired by the model proposed for VMR [14, 15, 16]. As shown on the right side of Figure 2, AM-DETR consists of a feature extractor and a DETR-based network the same as the VMR models. The key part of AM-DETR is the DETR-based network [14, 15, 16] that captures the temporal dependency between audio features, such as temporal changes in the audio features. Specifically, the architecture of AM-DETR is based on QD-DETR, which is simple but achieves high retrieval performance for VMR by capturing both the temporal dependency between video frames and cross-modal dependency between frames and text query [15].
@@
Finally, Transformer decoder transforms $\mathbf{z}^{\prime}$ to $K$ candidate audio moments $\hat{y}_{1},...,\hat{y}_{K}$ and their confidence scores $c_{1},...,c_{K}\in[0,1]$. The confidence score represents how plausible the corresponding predicted moment is. Here, $K$ is the hyperparameter controlling the number of outputs that assumes sufficiently larger than the number of the ground truth moment. To output $K$ candidates, the decoder also takes $K$ trainable anchors, which are parameters of the neural network controlling the number of output  [18] in addition to $\mathbf{z}^{\prime}$. Finally, the decoder is formulated as follows:
@@
Here, it is empirically known that using $c_{\phi(n)}$ not but $\log c_{\phi(n)}$ improves the performance in Eq (9[18]. We denote the optimal index as $\hat{\phi}=\underset{\phi\in\mathcal{P}}{\mathrm{argmin}}~\mathcal{L}_{\rm match}(\hat{y},y,c,\phi)$ where $\mathcal{P}$ is a set of every $K!$ permutation. Finally, the overall loss is obtained by the weighted sum of the moment loss and the CE loss as follows:
@@
Method
Feature
extractor
Contrastive
learning
Window
length
Clotho-Moment (eval) UnAV-100 subset TUT-Sound Events 2017
@@ V-A Model Configuration
Proposed model. To investigate the impact of the feature extractor, we used two feature extractors, CLAP [5] and VR [35]. VR is also a conventional audio retrieval model consisting of VAST [36] and RoBERTa [37]. To confirm whether pre-training of encoders based on contrast learning is effective in AMR, we also used VAST and RoBERTa without contrastive learning as the feature extractors. For comparison, we used the sliding window with 1, 4, and 7-second window lengths. The hyperparameter settings of the DETR-based network followed [15]. Both Transformer encoder and decoder consisted of two stacks of attention layers with 256 units and eight heads, and the dimension of the feed-forward layers was set to 1024. We used AdamW optimizer [38] with a learning rate of $1\times 10^{-4}$ and trained the models for 100 epochs with a batch size of 32.
Baseline. To measure how much AM-DETR trained with Clotho-Moment improves AMR performance by capturing time dependence, we prepared a baseline model using an existing speech retrieval model without training for AMR. The baseline model is a combination of the existing audio retrieval model [5] with the sliding window. To obtain audio moments, the baseline measures the similarities between audio and text embeddings for each window frame and then binarizes the similarities by comparing them to a threshold. The confidence score is defined as the average similarity within the moment. To improve performance, we applied a median filter to the resulting binary sequence. The threshold for the binarization and the length of median filters were tuned with the validation split of Clotho-Moment.
@@ V-C Results
Proposed vs Baseline. AM-DETR significantly outperformed the baseline in all metrics as shown in Table I (rows 4 and 6). Notably, the improvement of R1@0.7 and average mAP for UnAV-100 subset was 9.00 and 11.82 points. This result suggests that capturing the temporal dependency between audio features is important for AMR.
AMR performance on SED dataset In addition to the Clotho-Moment evaluation split and UnAV-100 subset, we also evaluate the AMR models on TUT sound events 2017 (TUT) [39] used in SED. Since the labels of TUT consist of too short words for the text query of AMR, we added several words with an average of 4.8 words to each label. Comparing the UnAV-100 subset, the performance on TUT was significantly low. This is because the ground-truth labels of TUT have very high and very high resolutions and contain many fine intervals. This property is different from that of Clotho-Moment, which contains few fine intervals.
Feature extractor. We investigate the impact of contrastive learning on the feature extractor. First, we found that contrastive learning of the feature extractor significantly improves the retrieval performance (rows 4 and 5). Notably, the improvement of R1@0.7 and average mAP for UnAV-100 subset was 29.00 and 23.11 points, respectively. Next, we investigate the impact of the architecture of the feature extractors. Comparing VR and CLAP (rows 5 and 6), the performance was comparable. Considering that VR outperformed CLAP in audio retrieval [35], this result indicates that an encoder superior in audio retrieval is not necessarily equally good in AMR.
@@
AMR as zero-shot SED. Because the AMR models can be seen as zero-shot SED as explained in Section II, we evaluated AM-DETR using the same metrics as for SED. In addition to AM-DETR, we evaluate a model that simply combines CLAP and the sliding window, which we call SED-CLAP. The thresholds for SED were tuned with the development set to maximize the F1 score, and the threshold was set to 0.5 if it was not tuned. The SED performance on TUT is shown in Table II and we found two insights. First, AM-DETR achieved the F1 score of over 41 points, which is 82% of the conventional SED model based on supervised learning [40], despite a zero-shot condition. Second, comparing SED-CLAP, AM-DETR is less sensitive to the threshold, which reduces tuning effort.
@@
TABLE II: Performance of zero-shot SED on TUT. All metrics are measured in micro average. The window lengths of SED-CLAP and AM-DETR were 7 and 1, respectively.